# Observational constraints on extended Proca-Nuevo gravity and cosmology

### v1

@@ Abstract
We confront massive Proca-Nuevo gravity with cosmological observations. The former is a non-linear theory involving a massive spin-1 field, that can be extended incorporating operators of the Generalized Proca class, and when coupled to gravity it can be covariantized in a way that exhibits consistent and ghost-free cosmological solutions, without experiencing instabilities and superluminalities at the perturbative level. When applied at a cosmological framework it induces extra terms in the Friedmann equations, however due to the special non-linear construction the field is eliminated in favor of the Hubble function. Thus, the resulting effective dark energy sector is dynamical, but with just one model parameter, namely the energy scale that controls the strength of the vector self-interactions. We use data from Supernovae Ia (SNIa) and Cosmic Chronometers (CC) observations and we construct the corresponding likelihood-contours for the free parameters. Interestingly enough, application of various information criteria, such as AIC, BIC and DIC, shows that the scenario of massive Proca-Nuevo gravity, although having exactly the same number of free parameters with $\Lambda$CDM concordance model, is more efficient in fitting the data. Finally, the reconstructed dark-energy equation-of-state parameter shows statistical compatibility with the model-independent, data-driven reconstructed one.
@@ 1 Introduction
The concordance model of cosmology, namely $\Lambda$CDM paradigm, which is based on general relativity with a comsological constant, on the standard-model particles, and on cold dark matter, has been proven very efficient in describing the Universe evolution both at the background and perturbation levels. Nevertheless, according to recent observations of various origins, $\Lambda$CDM predictions seem to be in tension with the data, as for instance the $H_{0}$ tension [1], the $\sigma_{8}$ tensions [2] etc (for reviews see [3, 4]). On the other hand, at the theoretical level, $\Lambda$CDM faces the cosmological constant problem [5, 6], while general relativity itself is non-renormalizable and thus it cannot be brought close to a quantum description [7, 8]. Hence, a large amount of research has been devoted in constructing gravitational modifications, namely theories that possess general relativity as a limit but presenting theoretical as well as phenomenological advantages (for a review see [9]).
A first subset of modified gravity theories emerges by enhancing the Einstein-Hilbert Lagrangian through the inclusion of supplementary terms. This results in a diverse array of formulations, encompassing $f(R)$ gravity [10, 11], $f(G)$ gravity [12], $f(P)$ gravity [13], and Lovelock gravity [14]. Nonetheless, an alternative trajectory involves transcending the conventional curvature-centric approach to gravity by integrating other geometrical parameters, such as torsion and non-metricity. More specifically, one can embark from the teleparallel equivalent of general relativity [15, 16], utilizing the torsion scalar $T$ as a Lagrangian and extending it to $f(T)$ gravity [17, 18]. Alternatively, one may extend the scope by adopting the symmetric teleparallel theory, utilizing the non-metricity scalar as the Lagrangian [19], and further extending it to $f(Q)$ gravity [20, 21]. All these classifications of gravitational theories demonstrate rich cosmological behaviors, captivating the keen interest of the scientific community [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48].
@@
Inspired by the structure of massive gravity, recently, an extended massive theory appeared in the literature, namely the massive Proca-Nuevo gravity [58]. It corresponds to a non-linear theory of a massive spin-1 field, that can be extended by incorporating operators of the Generalized Proca class without compromising the essential primary constraint crucial for consistency. When the theory is combined with gravity it can be covariantized in models that support consistent and ghost-free cosmological solutions. In particular, they exhibit hot Big Bang solutions featuring a late-time self-accelerating epoch, and additionally at the perturbative level they satisfy all stability and subluminality conditions, and intriguingly, gravitational waves propagate at the speed of light.
In the present work we desire to confront massive Proca-Nuevo gravity with cosmological data from Supernovae Ia (SNIa) and Cosmic Chronometers (CC) observations, and extract constraints on the model parameters. The plan of the work is the following: In Section 2 we present extended Proca-Nuevo gravity and we apply it at a cosmological framework. In Section 3 we provide the datasets that we will use in our analysis, and we analyze the various information criteria used for model comparison. Then, in Section 4 we perform the observational confrontation and we present the results. Finally, Section 5 is devoted to the Conclusions.
@@ 2 Extended Proca-Nuevo gravity and cosmology
Inserting all these in the first Friedmann equation (2.10) we obtain the simple algebraic equation
@@
where $E(z)\equiv H(z)/H_{0}$ is the normalized Hubble function. It is apparent that for large redshifts the term $E^{-2/3}$ becomes much less important, and the model effectively reduces to pure general relativity plus cold dark matter scenario. Moreover, note that by applying the first Friedmann equation at present time, the single model parameter, namely the scale $\Lambda$, has been eliminated in favor of $\Omega_{m0}$. Interestingly enough the scenario of extended Proca-Nuevo cosmology has exactly the same number of parameters with $\Lambda$CDM cosmology.
Lastly, by aid of equations (2.15),(2.16), we can define the effective equation-of-state parameter, namely $w_{DE}\equiv p_{DE}/\rho_{DE}$, as
@@ 3 Observational Data and Analysis
In this Section we review the datasets that we will employ in our analysis, namely data from Supernovae Ia (SNIa) observations and Cosmic Chronometers (CC) direct measurements of the Hubble function. Additionally, we provide the various information criteria used for model comparison.
3.1 Supernovae Ia (SNIa
@@
$\chi^{2}_{SNIa}\left(\phi^{\nu}_{\text{SNIa}}\right)={\bf\mu}_{\text{SNIa}}\,{\bf C}_{\text{SNIa},\text{cov}}^{-1}\,{\bf\mu}_{\text{SNIa}}^{T}\,,$ (3.1)
with ${\bf\mu}_{\text{\text{SNIa}}}=\{\mu_{1}-\mu_{\text{th}}(z_{1},\phi^{\nu})\,,\,...\,,\,\mu_{N}-\mu_{\text{th}}(z_{N},\phi^{\nu})\}$. The distance modulus is defined as $\mu_{i}=\mu_{B,i}-\mathcal{M}$, with $\mu_{B,i}$ the apparent magnitude at maximum in the rest frame for redshift $z_{i}$, and where $\mathcal{M}$ is a hyper-parameter representing uncertainties from various sources. Additionally, the theoretical form of the distance modulus is given by:
@@ 3.3 Likelihood analysis and model selection
The statistical vector has a dimension of $k$, comprising $\nu$ parameters of the model under consideration plus $\nu_{\text{hyp}}$ hyper-parameters from the utilized data sets, yielding $k=\nu+\nu_{\text{hyp}}$. In the above expression $\phi^{\mu}$ is the vector containing the constrained free parameters, which in our case is $\phi^{\mu}=\{\Omega_{m0},h,\mathcal{M}\}$ and $\mathcal{P}=\{CC,SNIa\}$. Lastly, to obtain the posterior distributions of the model parameters, given the data, we use a Markov Chain Monte Carlo (MCMC) sampler. Instead of the standard Metropolis - Hastings algorithm, to avoid the need of fine-tuning for the hyper-parameters, we use an affine-invariant Markov Chain Monte Carlo sampler as implemented within the open-source Python package emcee [66, 67], involving 1000 chains (walkers) and 2500 steps (states). Regarding the convergence of the MCMC algorithm, we use the traditional Gelman-Rubin criterion and also the auto-correlation time analysis.
@@ 5 Conclusions
When it is applied at a cosmological framework, massive Proca-Nuevo gravity induces extra terms in the Friedmann equations that can be collectively absorbed into an effective dark-energy sector. The interesting feature of the special non-linear construction is that the Klein-Gordon equation that arises form the scalar field parametrizing the vector ansatz, is non-dynamical, namely it is just a constraint imposing an algebraic relation between the field and the Hubble function. This allows to eliminate the field in favor of the Hubble function, and thus the effective dark energy sector finally depends only on the Hubble function, while the model has only one free parameter, namely the energy scale $\Lambda$ that controls the strength of the vector self-interactions. In summary, the cosmological scenario of massive Proca-Nuevo gravity is different than $\Lambda$CDM paradigm, but possessing exactly the same number of free parameters.

### v2

@@ Abstract
We confront massive Proca-Nuevo gravity with cosmological observations. The former is a non-linear theory involving a massive spin-1 field, that can be extended incorporating operators of the Generalized Proca class, and when coupled to gravity it can be covariantized in a way that exhibits consistent and ghost-free cosmological solutions, without experiencing instabilities and superluminalities at the perturbative level. When applied at a cosmological framework it induces extra terms in the Friedmann equations, however due to the special non-linear construction the field is eliminated in favor of the Hubble function. Thus, the resulting effective dark energy sector is dynamical, however it contains the same number of free parameters with the $\Lambda$CDM concordance model. We use data from Supernovae Ia (SNIa) and Cosmic Chronometers (CC) observations and we construct the corresponding likelihood-contours for the free parameters. Interestingly enough, application of various information criteria, such as AIC, BIC and DIC, shows that the scenario of massive Proca-Nuevo gravity, although having exactly the same number of free parameters with $\Lambda$CDM paradigm, it is more efficient in fitting the data. Finally, the reconstructed dark-energy equation-of-state parameter shows statistical compatibility with the model-independent, data-driven reconstructed one.
@@ 1 Introduction
The concordance model of cosmology, namely $\Lambda$CDM paradigm, which is based on general relativity with a cosmological constant, on the standard-model particles, and on cold dark matter, has been proven very efficient in describing the Universe evolution both at the background and perturbation levels. Nevertheless, according to recent observations of various origins, $\Lambda$CDM predictions seem to be in tension with the data, as for instance the $H_{0}$ tension [1], the $\sigma_{8}$ tensions [2] etc (for reviews see [3, 4]). On the other hand, at the theoretical level, $\Lambda$CDM faces the cosmological constant problem [5, 6], while general relativity itself is non-renormalizable and thus it cannot be brought close to a quantum description [7, 8]. Hence, a large amount of research has been devoted in constructing gravitational modifications, namely theories that possess general relativity as a limit but presenting theoretical as well as phenomenological advantages (for a review see [9]).
A first subset of modified gravity theories emerges by enhancing the Einstein-Hilbert Lagrangian through the inclusion of supplementary terms. This results in a diverse array of formulations, encompassing $f(R)$ gravity [10, 11], $f(G)$ gravity [12], $f(P)$ gravity [13], and Lovelock gravity [14]. Nonetheless, an alternative trajectory involves transcending the conventional curvature-centric approach to gravity by integrating other geometrical parameters, such as torsion and non-metricity. More specifically, one can embark from the teleparallel equivalent of general relativity [15, 16], utilizing the torsion scalar $T$ as a Lagrangian and extending it to $f(T)$ gravity [17, 18]. Alternatively, one may extend the scope by adopting the symmetric teleparallel theory, utilizing the non-metricity scalar as the Lagrangian [19], and further extending it to $f(Q)$ gravity [20, 21]. All these classifications of gravitational theories demonstrate rich cosmological behaviors, captivating the keen interest of the scientific community [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52].
@@
Inspired by the structure of massive gravity, the generalized Proca action for a vector field with derivative self-interactions, resulting in a theory with only three propagating degrees of freedom, was proposed in [62, 63]. In particular, a consistent, local theory of a massive vector field without the presence of ghost-like instabilities was constructed, and thus the cosmological solutions at both background and perturbation levels were studied [63]. The generalized Proca theory proves to have interesting cosmological phenomenology [64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79].
Recently, an extended Proca theory, namely Proca-Nuevo (PN) theory, appeared in the literature [80]. It corresponds to a non-linear theory of a massive spin-1 field, that can be extended by incorporating operators of the Generalized Proca class without compromising the essential primary constraint crucial for consistency. When the theory is combined with gravity it can be covariantized in models that support consistent and ghost-free cosmological solutions. In particular, they exhibit hot Big Bang solutions featuring a late-time self-accelerating epoch, and additionally at the perturbative level certain sub-classes of the theory satisfy all stability and subluminality conditions, and intriguingly, gravitational waves propagate at the speed of light. Moreover, further cosmological solutions of the theory have been studied in [81], while the complete analysis of the constraint algebra has been performed in [82, 83]. Finally, the quantum stability of Proca-Nuevo interactions was explored in [84], where it was found that Proca-Nuevo and generalized Proca theories have analogous behaviour at the quantum level, opening the door to speculations for being specific cases of a more general theory.
In the present work we desire to confront massive Proca-Nuevo gravity and cosmology with cosmological data from Supernovae Ia (SNIa) and Cosmic Chronometers (CC) observations, and extract constraints on the model parameters. The plan of the work is the following: In Section 2 we present extended Proca-Nuevo gravity and we apply it at a cosmological framework. In Section 3 we provide the datasets that we will use in our analysis, and we analyze the various information criteria used for model comparison. Then, in Section 4 we perform the observational confrontation and we present the results. Finally, Section 5 is devoted to the Conclusions.
@@ 2 Extended Proca-Nuevo gravity and cosmology
Thus, the first Friedmann equation (2.10) becomes $\Omega_{m}+\Omega_{DE}=1$, which applied at present time, i.e. at $z=0$, and using (2.15), gives
$\frac{\Lambda^{16/3}c_{m}y^{2/3}}{6M_{Pl}^{8/3}H_{0}^{8/3}}=1-\Omega_{m0}.$ (2.19)
Note that although the scenario has two intrinsic parameters, namely $\Lambda$ and $m$, the specific combination in which both of them appear can be eliminated in favour of just $\Omega_{m0}$. Inserting (2.19) back into (2.15) we can eliminate $c_{m}y^{2/3}$, as well as $\Lambda$, and therefore we obtain
$\rho_{DE}=3(1-\Omega_{m0})M_{Pl}^{2}H_{0}^{8/3}H^{-2/3}.$ (2.20)
Interestingly enough, in the scenario at hand the effective dark energy density depends only on the present matter density parameter $\Omega_{m0}$, as well as on the usual normalization factor, i.e. the present Hubble parameter $H_{0}$, and thus it has the same number of free parameters as $\Lambda$CDM cosmology, which for comparison we remind that it has
$\rho_{DE}|_{{}_{\Lambda CDM}}=3(1-\Omega_{m0})M_{Pl}^{2}H_{0}^{2}.$ (2.21)
In summary, the scenario of extended Proca-Nuevo cosmology, assuming a spatially flat universe and considering only pressureless matter, has exactly the same number of parameters with $\Lambda$CDM cosmology, however its cosmological behavior will in general be different.
Inserting all these in the first Friedmann equation (2.10), we obtain the simple algebraic equation
@@
where $E(z)\equiv H(z)/H_{0}$ is the normalized Hubble function. Furthermore, by aid of equations (2.15),(2.16), we can define the effective equation-of-state parameter, namely $w_{DE}\equiv p_{DE}/\rho_{DE}$, as
@@
In summary, it is apparent that for large redshifts the term $E(z)^{-2/3}$ becomes much less important, and the model effectively reduces to pure general relativity plus cold dark matter scenario. On the other hand, at late times, where $H\rightarrow H_{0}$ and thus $E(z)\rightarrow 1$, we observe that $\rho_{DE}$ behaves close to $\rho_{DE}|_{{}_{\Lambda CDM}}$ and for $z=0$ we obtain full coincidence. Hence, the scenario at hand of Proca-Nuevo cosmology has exactly the same number of parameters with $\Lambda$CDM cosmology, it reduces to the latter at early and present times, however at intermediate times it exhibits a non-trivial deviation. In the next section we desire to see whether this intermediate-time deviation can improve the fitting with the data comparing to $\Lambda$CDM paradigm.
@@ 3 Observational Data and Analysis
In this Section we review the datasets that we will employ in our analysis, namely data from Supernovae Ia (SNIa) observations and Cosmic Chronometers (CC) direct measurements of the Hubble function. Additionally, we provide the various information criteria used for model comparison. In what follows, $\phi^{\lambda}$ is the statistical vector, that contains the free parameters at hand.
3.1 Supernovae Ia (SNIa)
@@
$\chi^{2}_{SNIa}\left(\phi^{\nu}\right)={\bf\mu}_{\text{SNIa}}\,{\bf C}_{\text{SNIa},\text{cov}}^{-1}\,{\bf\mu}_{\text{SNIa}}^{T}\,,$ (3.1)
with ${\bf\mu}_{\text{\text{SNIa}}}=\{\mu_{1}-\mu_{\text{th}}(z_{1},\phi^{\nu})\,,\,...\,,\,\mu_{N}-\mu_{\text{th}}(z_{N},\phi^{\nu})\}$. The distance modulus is defined as $\mu_{i}=\mu_{B,i}-\mathcal{M}$, with $\mu_{B,i}$ the apparent magnitude at maximum in the rest frame for redshift $z_{i}$. The introduction of the free parameter $\mathcal{M}$ is essential due to the dependence of the observable distance modulus, $\mu_{obs}$, on assumptions related to $H_{0}$ and the fiducial cosmology. By absorbing artifacts originating from the transformation of flux observations to distance modulus estimates, $\mathcal{M}$ enables independence from the fiducial cosmology, as discussed in [87]. Additionally, the theoretical form of the distance modulus is given by:
@@ 3.3 Likelihood analysis and model selection
The statistical vector has a dimension of $k$, comprising $\nu$ parameters of the model under consideration plus $\nu_{\text{hyp}}$ hyper-parameters from the utilized data sets, yielding $k=\nu+\nu_{\text{hyp}}$. In the above expression $\phi^{\mu}$ is the vector containing the free parameters, which in our case is $\phi^{\mu}=\{\Omega_{m0},h,\mathcal{M}\}$ and $\mathcal{P}=\{CC,SNIa\}$. Lastly, to obtain the posterior distributions of the model parameters, given the data, we use a Markov Chain Monte Carlo (MCMC) sampler. Instead of the standard Metropolis - Hastings algorithm, to avoid the need of fine-tuning for the hyper-parameters, we use an affine-invariant Markov Chain Monte Carlo sampler as implemented within the open-source Python package emcee [92, 93], involving 1000 chains (walkers) and 2500 steps (states). Regarding the convergence of the MCMC algorithm, we use the traditional Gelman-Rubin criterion and also the auto-correlation time analysis.
@@ 5 Conclusions
When it is applied at a cosmological framework, massive Proca-Nuevo gravity induces extra terms in the Friedmann equations that can be collectively absorbed into an effective dark-energy sector. The interesting feature of the special non-linear construction is that the Klein-Gordon equation that arises form the scalar field parametrizing the vector ansatz, is non-dynamical, namely it is just a constraint imposing an algebraic relation between the field and the Hubble function. This allows to eliminate the field in favor of the Hubble function, and thus the effective dark energy sector finally depends only on the Hubble function. In summary, the cosmological scenario of massive Proca-Nuevo gravity is different than $\Lambda$CDM paradigm, but possessing exactly the same number of free parameters.